# How Worst-Case Are Adversarial Attacks? Linking Adversarial and Statistical Robustness

### v1

@@ Abstract
Adversarial attacks are widely used to evaluate model robustness, yet their validity as proxies for robustness to random perturbations remains debated. We ask whether an adversarial perturbation provides a representative estimate of robustness under random noise of the same magnitude, or instead reflects an atypical worst-case event. To this end, we introduce a probabilistic metric that quantifies noisy risk with respect to directionally biased perturbation distributions, parameterized by a concentration factor $\kappa$ that interpolates between isotropic noise and adversarial direction. Using this framework, we study the limits of adversarial perturbations as estimators of noisy risk by proposing an attack strategy designed to operate in regimes statistically closer to uniform noise. Experiments on ImageNet and CIFAR-10 systematically benchmark widely used attacks, highlighting when adversarial success meaningfully reflects noisy risk and when it fails, thereby informing their use in safety-oriented evaluation.
@@ 1 Introduction
Adversarial examples, defined as $\varepsilon$-bounded perturbations intentionally crafted to alter a model’s prediction, have been extensively studied in the security literature (Biggio & Roli, 2018), where it is typically assumed that an attacker can deliberately manipulate digital inputs. Beyond security, adversarial examples have also been widely adopted by the safety community as a tool to evaluate and characterize worst-case robustness to perturbations (Huang et al., 2020). Under this perspective, the success of an adversarial attack is often treated as a conservative indicator of safety, as a sample is deemed non-robust whenever an attack succeeds.
@@
Despite their widespread adoption, current models often exhibit strong robustness to random noise, while they are quite vulnerable to adversarial attacks (Carlini et al., 2019). This discrepancy raises fundamental questions about whether adversarial attacks truly reflect robustness to perturbations that are likely to be encountered in practice.
@@
Motivating the gap. Figure 1 illustrates the discrepancy: for clean samples that are correctly classified, we report misclassification rates obtained by applying adversarial attacks and by a Monte Carlo (MC) error search based on $T$ uniformly sampled perturbations of the same $\varepsilon$ magnitude, where a sample is deemed misclassified if at least one of the $T$ perturbations induces an incorrect prediction. Even for very large $T$, misclassification rate under random perturbations remains substantially lower than that induced by adversarial attacks, becoming comparable only at large $\varepsilon$.  [^1] T=700, [^0] T=700{,}000 corresponds a probability on the order of $10^{-6}$, comparable to extremely rare catastrophic events (Plait, 2008).
@@
In the literature, several analyses have questioned the practical relevance of adversarial attacks for safety evaluation (Carlini et al., 2019; Fawzi et al., 2016). While insightful, they may be interpreted as discouraging the use of attacks altogether, rather than clarifying when they remain informative. This motivates tools that explicitly relate adversarial success to noisy risk, understood as model vulnerability to unintentional perturbations of similar nature.
@@
This work introduces a probabilistic measure of noisy risk associated with an adversarial direction $v$ produced by an attack. We evaluate it by assessing robustness over a neighborhood of directions with fixed magnitude around $v$, modeled by a Gaussian distribution whose mean is aligned with $v$ and whose concentration is controlled by a parameter $\kappa$. Low values of $\kappa$ yield distributions close to uniform noise on the $\ell_{p}$-sphere, whereas larger values increasingly concentrate mass around $v$. Consequently, high attack success at low $\kappa$ indicates statistically plausible vulnerabilities, while success only at high $\kappa$ reflects localized corner cases.
@@
Building on this analysis, we then introduce a directional noisy attack that explicitly seeks adversarial perturbations maximizing attack success in low-$\kappa$ regimes, thereby better capturing more statistically plausible noisy risk regions. The attack injects controlled noise during the optimization process, sampled from a distribution aligned with the adversarial direction available at each attack iteration.
@@
  • a definition and analysis of directional noisy risk, providing a probabilistic understanding for relating adversarial perturbations to stochastic noise of the same magnitude;
  • a directional noisy attack that optimizes adversarial success in low-$\kappa$ regimes, probing the limits of adversarial attacks as proxies for perturbation-based robustness;
  • a systematic experimental evaluation across multiple settings, attacks and models, offering practical insights and guidelines for robustness evaluations.
@@ 2 Related work
Overall, there remains substantial ambiguity regarding how adversarial attacks should be treated for safety assessment. (Olivier & Raj, 2023) studies the geometric abundance of adversarial perturbations via constrained directional searches, while (Guo et al., 2025) analyzes how adversarial behavior evolves as the perturbation magnitude increases. In contrast, this work aims to provide a statistical understanding of a given adversarial direction at a fixed magnitude, thereby offering a tool for benchmarking attacks and for relating adversarial behavior to statistical noise of the same nature.
@@ 3 Background and Problem Formulation
Let $f:\mathcal{X}\to\mathcal{Y}$ be a neural network. For semplicity, we consider the case of image classifier, where $\mathcal{Y}$ is the output space and denote by $\hat{y}(x)=\arg\max_{c\in\mathcal{Y}}f_{c}(x)$ the predicted label. Let $x\in\mathcal{X}$ be a clean input with ground-truth label $y\in\mathcal{Y}$. We introduce two notions of robustness and examine the challenge of bridging them.
@@
This formulation relies on the existence of a single adversarial perturbation $\delta$, given by an attack $\mathcal{A}$, to deem a point non-robust, without modeling the probability of such a perturbation under a statistical noise perspective.
Noisy-bounded robustness. From a general perspective, robustness to noise is a large concept that can be formalized with respect to selected corruptions and transformations (Hendrycks & Dietterich, 2019; Carvalho et al., 2025). To establish a direct correspondence with adversarial perturbations, we instead focus on stochastic perturbations of fixed magnitude. We consider perturbations sampled uniformly from the $\ell_{p}$-sphere of radius $r$ centered at $x$. Such a distribution can be approximated [^2] Exact for $\ell_{2}$; approximate for $\ell_{p}$, $p\neq 2$ (see Appendix A). via a Gaussian by $\eta=\Pi_{\|\cdot\|_{p}=r}(\xi),$ where $\xi\sim\mathcal{N}(0,I),$ and $\Pi_{\|\cdot\|_{p}=r}$ denotes projection onto the $\ell_{p}$-sphere of radius $r$. Under this construction, the noisy risk error of a correctly classified sample ($x$ such that $\hat{y}(x)=y$) is defined as
@@
While noisy error provides a statistically meaningful notion for safety assessment, exhaustively estimating it in practice via Monte Carlo sampling may require prohibitive evaluation time (Cohen et al., 2019). Consequently, adversarial attacks are treated as pointwise robustness surrogates, even though they capture fundamentally different failure notions: the existence of an adversarial perturbation $\delta$ only certifies the presence of a specific error direction and does not imply that such a direction is representative under stochastic perturbations of the same magnitude (Eq. 2 for $r=\|\delta\|_{p}$). As a result, the common empirical practice of inferring noisy risk behavior from adversarial success (Eq. 1) is generally unjustified, as any potential relationship critically depends on the properties of the attack $\mathcal{A}$ and on the geometry of the model loss landscape (Li et al., 2018). This motivates the need for tools that quantify how adversarial directions relate to statistically plausible perturbations. [^3] We use the same $p$-norm for both adversarial attacks and noisy risk analysis; this choice is not restrictive (see Appendix A).
4 Directional Noisy Risk Metric
We study whether the unit adversarial direction $v\;=\;\frac{\delta}{\|\delta\|_{2}}$ points toward a representative noisy failure region on a $\ell_{p}$-sphere of a radius $r$, thereby linking adversarial robustness with noisy risk analysis. To explore this, rather than focusing on isolated perturbations or uniform ones, we aim to characterize whether misclassification persists under stochastic perturbations that are biased toward $v$. Specifically, we generate perturbations by sampling from a Gaussian distribution whose mean is shifted along the direction $v$ and then projecting onto the $r$-sphere:
@@ 4 Directional Noisy Risk Metric
where $\kappa\geq 0$ controls the strength of the directional bias: $\kappa=0$, this reduces to uniform noise on the sphere (as in Eq.2), while larger values of $\kappa$ progressively concentrate probability mass around the direction $v$. We thus define the directional noisy risk error of a direction $v$ as
@@
In practice, $\mathcal{R}$ can be estimated numerically as the misclassification rate over $n$ sampled perturbations. For direct comparison with an adversarial perturbation $\delta$, we consider $r=\|\delta\|_{p}$. By varying $\kappa$ as an evaluation parameter, we can analyze different levels of noisy risk, ranging from uniform noise on the $\ell_{p}$-sphere to perturbations concentrated around the adversarial direction $v$ as $\kappa\to\infty$, which progressively resemble the original attack success rate.
Figure 2: Illustration of the proposed metric in a 2D-example. (a) and (b) compare two adversarial directions for the same input $x_{0}$ that differ in their directional noisy risk values $\mathcal{R}_{\kappa=5}$. (c) depicts a corner-case adversarial direction with a low noisy risk score.
@@
Figure 2 illustrates the proposed metric in a simplified two-dimensional setting. Fig. 2(a) and  Fig.2(b) show two adversarial examples on the same $\ell_{2}$-sphere centered at $x_{0}$ that induce misclassification but differ in their directional noisy risk error $\mathcal{R}$. Although both perturbations are adversarial, the example in Fig. 2(b) yields a much higher $\mathcal{R}$ (here with $\kappa=5$), indicating that a larger neighborhood of nearby perturbations remains misclassified. In contrast, the adversarial direction in Fig. 2(a) corresponds to a more localized failure region, resulting in a lower noisy risk score. Figure 2(c) considers a different input point $x_{1}$, where misclassification occurs only along a narrow, corner region of the decision boundary. Despite being adversarial, this perturbation yields a very low noisy risk for $\kappa=5$, revealing its highly localized and non-representative nature.
@@
From a practical perspective, to match a noisy risk scenario, $\kappa$ should be chosen such that the resulting perturbations remain statistically close to uniform noise on the $\ell_{p}$-sphere. For $p=2$, consider $D$ the input dimensions and write $\xi=\kappa v+z$, with $z\sim\mathcal{N}(0,I)$, by concentration of measure, the norm of the random component satisfies $\|z\|_{2}\approx\sqrt{D}$ with high probability. As a consequence, the directional shift $\kappa v$ remains negligible compared to the $z$ as long as $\kappa\ll\sqrt{D}$, while for $\kappa\gtrsim\sqrt{D}$ the perturbations increasingly concentrate around $v$. However, although $\sqrt{D}$ emerges as a natural scale from a norm-based perspective, in practice, values of $\kappa$ on the order of $\sqrt{D}$ already induce a strongly directional perturbation distribution, due to the rapid concentration of probability mass around the mean direction in high-dimensional spaces. As a result, we propose $\kappa^{*}=D^{1/4}$ as an empirical yet more conservative reference point. This choice lies well within a stochastic regime, while still introducing a measurable directional preference after projection. We further justify this choice experimentally in Sec. 6.3 for both $p=2$ and $p=\infty$ perturbations. Additional analyses are in Appendix B.
@@ 5 Directional Noisy Attack
20:  $g\leftarrow\mu g+\mathrm{Normalize_{p}}(\bar{g})$ #momentum
@@
Well-known adversarial strategies, such as PGD (Madry et al., 2018), DeepFool (Moosavi-Dezfooli et al., 2016), and CW (Carlini & Wagner, 2017), iteratively follow the local steepest-ascent direction of the loss to maximize a task loss $\mathcal{L}$ in untargeted settings, i.e. $\max_{\lVert\delta\rVert_{p}\leq\varepsilon}\mathcal{L}\!\left(f(x+\delta),y\right).$
While effective at identifying worst-case perturbations, these methods provide no probabilistic characterization of whether the resulting adversarial direction corresponds to a broad failure region or merely an isolated decision-boundary crossing (Fawzi et al., 2016). To favor adversarial directions associated with high-probability failure regions, we follow an alternative objective that directly maximizes the proposed directional noisy risk error at a selected concentration ${\kappa}_{\text{adv}}$
@@
$\bar{g}^{t}\;\approx\;\mathbb{E}_{\xi\sim\mathcal{N}(\kappa_{\text{adv}}v^{t},I)}\Bigl[\nabla_{x}\ell\bigl(f(x_{0}+\eta),y\bigr)\Bigr],$ (6)
@@
Table 1: Directional noisy-risk errors under $\ell_{\infty}$ attacks ($\varepsilon\in\{4/255,8/255\}$) and $\ell_{2}$ attacks ($\varepsilon\in\{1.0,2.0\}$). $\bar{\mathcal{R}}_{\kappa^{*}}$ denotes the average directoinal noisy-risk error on the ImageNet-Val evaluated at $\kappa^{*}=D^{1/4}$, while $\kappa_{0.25}$ denotes the smallest concentration parameter $\kappa$ estimated for which $\bar{\mathcal{R}}\geq 0.25$. ASR denotes the standard attack success rate, reported for comparison.

### v2

# How Worst-Case Are Adversarial Attacks? Linking Adversarial and Perturbation Robustness

@@ Abstract
Adversarial attacks are widely used to identify model vulnerabilities; however, their validity as proxies for robustness to random perturbations remains debated. We ask whether an adversarial example provides a representative estimate of misprediction risk under stochastic perturbations of the same magnitude, or instead reflects an atypical worst-case event. To address this question, we introduce a probabilistic analysis that quantifies this risk with respect to directionally biased perturbation distributions, parameterized by a concentration factor $\kappa$ that interpolates between isotropic noise and adversarial directions. Building on this, we study the limits of this connection by proposing an attack strategy designed to probe vulnerabilities in regimes that are statistically closer to uniform noise. Experiments on ImageNet and CIFAR-10 systematically benchmark multiple attacks, revealing when adversarial success meaningfully reflects robustness to perturbations and when it does not, thereby informing their use in safety-oriented robustness evaluation.
@@ 1 Introduction
Adversarial examples, defined as bounded perturbations intentionally crafted to alter a model prediction, have been extensively studied in the security literature (Biggio & Roli, 2018), where it is typically assumed that an attacker can deliberately manipulate digital inputs. Beyond security, adversarial examples have also been widely adopted by the safety community as a tool to evaluate and characterize worst-case robustness to perturbations (Huang et al., 2020). From this perspective, the success of an adversarial attack is often treated as a conservative indicator of safety, as a sample is deemed non-robust whenever an attack succeeds.
@@
Despite their widespread adoption, current models often exhibit strong robustness to random noise, while remaining highly vulnerable to adversarial attacks (Carlini et al., 2019). This discrepancy raises questions about whether adversarial attacks reflect robustness to statistically likely perturbations.
@@
Motivating the gap. Figure 1 illustrates this discrepancy: for clean samples that are correctly classified, we report misclassification rates obtained by applying adversarial attacks and a Monte Carlo (MC) error search based on $T$ uniformly sampled perturbations of a fixed $\varepsilon$ magnitude, where a sample is deemed misclassified if at least one of the $T$ perturbations induces an incorrect prediction. Even for very large $T$, misclassification rate under $\varepsilon$-random perturbations remains substantially lower than that induced by adversarial attacks, becoming comparable only at large $\varepsilon$.  [^1] T=700, [^0] T=700{,}000 corresponds a probability on the order of $10^{-6}$, comparable to extremely rare catastrophic events (Plait, 2008).
@@
In the literature, several analyses have questioned the practical relevance of adversarial attacks for safety evaluation (Carlini et al., 2019; Fawzi et al., 2016). While insightful, these works may discourage the use of attacks altogether, rather than clarifying when they remain informative. This motivates tools that explicitly relate adversarial success to perturbation risk, defined as model vulnerability to stochastic perturbations of comparable magnitude.
@@
This work introduces a probabilistic analysis of perturbation risk associated with an adversarial direction $v$ produced by an attack. We evaluate it by assessing robustness over a neighborhood of directions with fixed magnitude around $v$, modeled by a Gaussian distribution whose mean is aligned with $v$ and whose concentration is controlled by a parameter $\kappa$. Low values of $\kappa$ yield distributions close to uniform noise on the $\ell_{p}$-sphere, whereas larger values increasingly concentrate mass around $v$. Consequently, high attack success at low $\kappa$ indicates statistically plausible vulnerabilities, while success only at high $\kappa$ reflects localized corner cases.
@@
Building on this analysis, we then introduce a directional noisy attack that explicitly seeks adversarial perturbations maximizing attack success in low-$\kappa$ regimes, thereby better capturing stochastic failure regions. The attack injects controlled noise during optimization, sampled from a distribution aligned with the current adversarial direction and parameterized by a user-defined concentration.
@@
  • a definition of directional perturbation risk, providing a probabilistic understanding for relating adversarial examples to stochastic perturbations of the same magnitude;
  • a directional noisy attack optimizing adversarial success in low-$\kappa$ regimes, probing the limits of adversarial attacks as proxies for stochastic robustness;
  • a systematic experimental evaluation across multiple settings, attacks, and models, providing practical guidance for safety-oriented robustness assessment.
@@ 2 Related work
Overall, there remains substantial ambiguity regarding how adversarial attacks should be treated for safety assessment. (Olivier & Raj, 2023) studied the geometric abundance of adversarial perturbations via constrained directional searches, while (Guo et al., 2025) analyzed how adversarial behavior evolves as the perturbation magnitude increases. (Rice et al., 2021) proposed a continuum of robustness notions through functional loss norms. In contrast, this work provides a probabilistic characterization of a given adversarial direction, offering a tool for relating adversarial behavior to stochastic perturbations of the same magnitude.
@@ 3 Background and Problem Formulation
Let $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{Y}|}$ be a neural network used for classification, where $\mathcal{Y}$ denotes the set of class labels. We define $\hat{y}(x)=\arg\max_{c\in\mathcal{Y}}f_{c}(x)$ as the predicted label. Let $x\in\mathcal{X}$ be a clean input with ground-truth label $y\in\mathcal{Y}$. We introduce two notions of robustness and examine the challenge of bridging them.
@@
This formulation relies on the existence of an adversarial perturbation $\delta$, given by an attack $\mathcal{A}$, to deem a point non-robust, without modeling the probability of such a perturbation under a statistical noise perspective.
Robustness to fixed-magnitude perturbations. From a general perspective, robustness to noise is a broad concept that can be formalized with respect to selected corruptions and transformations (Hendrycks & Dietterich, 2019; Carvalho et al., 2025). To establish a direct correspondence with adversarial examples, we focus on stochastic perturbations of fixed magnitude. We consider perturbations sampled uniformly from the $\ell_{p}$-sphere of radius $r$ centered at $x$. Such a distribution can be approximated [^2] Exact for $\ell_{2}$; approximate for $\ell_{p}$, $p\neq 2$ (see Appendix A). via a Gaussian by $\eta=\Pi_{\|\cdot\|_{p}=r}(\xi),$ where $\xi\sim\mathcal{N}(0,I),$ and $\Pi_{\|\cdot\|_{p}=r}$ denotes projection onto the $\ell_{p}$-sphere of radius $r$. Under this construction, the perturbation risk of a correctly classified sample $x$ (i.e., $\hat{y}(x)=y$) is defined as
@@
While perturbation risk provides a statistically meaningful notion for safety assessment, exhaustively estimating it in practice via Monte Carlo sampling may require prohibitive evaluation time (Cohen et al., 2019). Consequently, adversarial attacks are treated as pointwise robustness surrogates, even though they capture fundamentally different failure notions: the existence of an adversarial perturbation $\delta$ only certifies the presence of a specific error direction and does not imply that such a direction is representative under stochastic perturbations of the same magnitude (Eq. 2 for $r=\|\delta\|_{p}$). As a result, the common empirical practice of inferring perturbation risk behavior from adversarial success (Eq. 1) is generally unjustified, as any potential relationship critically depends on the properties of the attack $\mathcal{A}$ and on the geometry of the model loss landscape (Li et al., 2018). This motivates the need for tools that quantify how adversarial directions relate to statistically plausible perturbations. [^3] We use the same $p$-norm for both adversarial attacks and perturbation risk; this choice is not restrictive (see Appendix A).
4 Directional Perturbation Risk Metric
We study whether the unit adversarial direction $v\;=\;\frac{\delta}{\|\delta\|_{2}}$ points toward a representative noisy failure region on a $\ell_{p}$-sphere of a radius $r$, thereby linking adversarial and perturbation robustness. To explore this, rather than focusing on isolated perturbations or uniform ones, we aim to characterize whether misclassification persists under stochastic perturbations that are biased toward $v$. Specifically, we generate perturbations by sampling from a Gaussian distribution whose mean is shifted along the direction $v$ and then projecting onto the $r$-sphere:
@@ 4 Directional Perturbation Risk Metric
where $\kappa\geq 0$ controls the strength of the directional bias: $\kappa=0$, this reduces to uniform noise on the sphere (as in Eq.2), while larger values of $\kappa$ progressively concentrate probability mass around the direction $v$. We thus define the directional perturbation risk of a direction $v$ as
@@
In practice, $\mathcal{R}$ can be estimated numerically as the misclassification rate over $n$ samples. For direct comparison with an adversarial perturbation $\delta$, we consider $r=\|\delta\|_{p}$. By varying $\kappa$ as an evaluation parameter, we can analyze different levels of perturbation risk, ranging from uniform noise on the $\ell_{p}$-sphere to perturbations concentrated around the adversarial direction $v$ as $\kappa\to\infty$, which progressively resemble the original attack success rate.
Figure 2: Illustrative examples. (a) and (b) compare two adversarial directions for the same input $x_{0}$ that differ in their directional perturbation risk, computed with $\kappa=5$. (c) depicts a corner-case adversarial direction with low perturbation risk.
@@
Figure 2 illustrates the proposed metric in a simplified two-dimensional setting. Fig. 2(a) and  Fig.2(b) show two adversarial examples on the same $\ell_{2}$-sphere centered at $x_{0}$ that induce misclassification but differ in their directional perturbation risk $\mathcal{R}$. Although both attacks are successful, the example in Fig. 2(b) yields a much higher $\mathcal{R}$ (here with $\kappa=5$), indicating that a larger neighborhood of nearby perturbations remains misclassified. In contrast, the adversarial direction in Fig. 2(a) corresponds to a more localized failure region, representing a lower perturbation risk. Figure 2(c) considers a different input point $x_{1}$, where misclassification occurs only along a narrow, corner region of the decision boundary. Despite being adversarial, this example yields a very low perturbation risk for $\kappa=5$, revealing its highly localized and non-representative nature.
@@
From a practical perspective, to match a stochastic scenario, $\kappa$ should be chosen such that the resulting perturbations remain statistically close to uniform noise on the $\ell_{p}$-sphere. For $p=2$, consider $D$ the input dimensions and write $\xi=\kappa v+z$, with $z\sim\mathcal{N}(0,I)$, by concentration of measure, the norm of the random component satisfies $\|z\|_{2}\approx\sqrt{D}$ with high probability. As a consequence, the directional shift $\kappa v$ remains negligible compared to the $z$ as long as $\kappa\ll\sqrt{D}$, while for $\kappa\gtrsim\sqrt{D}$ the perturbations increasingly concentrate around $v$. However, although $\sqrt{D}$ emerges as a natural scale from a norm-based perspective, in practice, values of $\kappa$ on the order of $\sqrt{D}$ already induce a strongly directional perturbation distribution, due to the rapid concentration of probability mass around the mean direction in high-dimensional spaces. As a result, we propose $\kappa^{*}=D^{1/4}$ as an empirical yet more conservative reference point. This choice lies well within a stochastic regime, while still introducing a measurable directional preference after projection. We further justify this choice experimentally in Section 6.3 for both $p=2$ and $p=\infty$ perturbations. Additional analyses are in Appendix B.
@@ 5 Directional Noisy Attack
20:  $g\leftarrow\mu g+\mathrm{Normalize_{p}}(\bar{g})$
@@
Well-known adversarial strategies, such as PGD (Madry et al., 2018), DeepFool (Moosavi-Dezfooli et al., 2016), and CW (Carlini & Wagner, 2017), iteratively follow the local steepest-ascent direction of the task objective to maximize $\mathcal{L}$ in untargeted settings, i.e $\max_{\lVert\delta\rVert_{p}\leq\varepsilon}\mathcal{L}\!\left(f(x+\delta),y\right).$
While effective at identifying worst-case perturbations, these methods provide no probabilistic characterization of whether the resulting adversarial direction corresponds to a broad failure region or merely an isolated point (Fawzi et al., 2016). To favor adversarial directions associated with high-probability failure regions, we follow an alternative objective that directly maximizes the proposed directional perturbation risk at a selected concentration ${\kappa}_{\text{adv}}$
@@
$\bar{g}^{t}\;\approx\;\mathbb{E}_{\xi\sim\mathcal{N}(\kappa_{\text{adv}}v^{t},I)}\Bigl[\nabla_{x}\ell\bigl(f(x+\eta),y\bigr)\Bigr],$ (6)
@@
Table 1: Directional perturbation risk under $\ell_{\infty}$ attacks ($\varepsilon\in\{4/255,8/255\}$) and $\ell_{2}$ attacks ($\varepsilon\in\{1.0,2.0\}$). $\bar{\mathcal{R}}_{\kappa^{*}}$ denotes the average directional perturbation risk on the ImageNet-Val evaluated at $\kappa^{*}=D^{1/4}$, while $\kappa_{0.25}$ denotes the smallest concentration parameter $\kappa$ estimated for which $\bar{\mathcal{R}}\geq 0.25$. ASR denotes the standard attack success rate, reported for comparison.